A stock is expected to pay a dividend of $0.75 at the end of the year (i.e., D1 = $0.75), and it should continue to grow at a constant rate of 3% a year. If its required return is 15%, what is the stock's expected price 1 year from today? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Stock price in 1 year (P1) D2=(r-g) Here, =0.75*(1+3%) Expected dividend (D2) Required return (r) Growth rate (g) 0.77 15% 3%

*Please rate thumbs up